Tom sold stock valued at $300,000. Its basis in the stock was $250,000. For the stock, Tom received $150,000 in cash in the current year and a note providing Tom with $150,000 in the subsequent year. Assuming Tom wants to defer as much tax as possible, what is Tom recognized gain in the current and subsequent year, respectively?

  1. $0; $50,000.
  2. $10,000; $40,000.
  3. $25,000; $25,000.
  4. $50,000; $0.
  5. None of the Above.
Answer #1

Answer: Option (c) $25,000; $25,000

Gross profit on sale = (300,000-250,000)/300,000 = 16.67%

Gain recognized in

Year 1 = 150,000 * 16.67% = 25,000

Year 2 = 150,000 * 16.67% = 25,000